8, ?, 64, 125, 216
1,2,9,3,6,3,1,?
1,3,4,7,11, ?
84178, 41788, 17884, 78841, ?


Answer by addingup(3677)

8, ?, 64, 125, 216 Missing number: 27. Why? These are CUBE NUMBERS, as in x^3
1*1*1 = 1
2*2*2 = 8
3*3*3 = 27
4*4*4 = 64
5*5*5 = 125
6*6*6 = 216
`````````````````````````````````````````````
Next: The next number is the sum of the previous two. After that, the next number is the sum of the previous two:
1,3,4,7,11, ?
1+3=4
3+4=7
4+7=11
7+11=18
`````````````````````````````````````````````
In this sequence, you move the first digit to the end to get the next number:
84178 41788 17884 78841 --> 88417
84178 take the first number (8) and move it to the back, you get:
41788 Now take the 4 and move it to the back:
17884 And the 1:
78841 And the 7, and it gives you the answer:
88417 If you do it one more time you would be back at the beginning: 84178
Good luck.
John